OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies inversely as x and y=6 and x=9

OpenStudy (anonymous):

y=K*1/x

OpenStudy (anonymous):

6=k*1/9 k=54

OpenStudy (anonymous):

The form for inverse variation is y=k/x where k is the constant of variation. Just sub for x and y to get the comnstant: 6=k/9 or k=54, thus the equation is y=54/x

OpenStudy (anonymous):

y=54/x